OpenStudy (anonymous):

Convert each measure to radians and state answer in terms of pi. a.) 40(degrees) b.) -570(degrees)

jimthompson5910 (jim_thompson5910):

I'll do the first one to get you started a) 40 degrees multiply the given degree measurement by pi/180 to convert to radian mode 40*(pi/180) = 40pi/180 = (40/180)pi = (2/9)pi = 2pi/9 so \[\Large 40^{\circ} = \frac{2\pi}{9} \text{radians}\]

OpenStudy (anonymous):

ok idk where half that came from? can you break it down so i can do the next one?

jimthompson5910 (jim_thompson5910):

which steps are throwing you off?

jimthompson5910 (jim_thompson5910):

Let's do another example say 15 degrees

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

multiply by the fraction pi/180